Solve the system of equation by graphing
y=-2x+2
7x+y=7
If you could please show all working and produce the end result on a graph, it will be greatly appreciated. Thank you.

Answer by ewatrrr(24785)


Hi
y=-2x+2
7x+y=7 OR y = -7x + 7
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept to graph the lines:
y = -2x + 2 |slope m = -2, y-intercept is Pt (0,2) |Pt(1,0) another point
y = -7x + 7 |slope m = -7, y-intercept is Pt (0,7) |Pt(1,0) anther point
Plotting the y-intercepts and using their slopes to find another point on each line
connecting these points concludes the graphing
Note: these Lines have Pt(1,0) in common